Calculate the molarity of the hydrochloric acid.
1
Expert's answer
2020-05-26T14:21:08-0400

Hydrochloric acid is a saturated aqueous solution of hydrogen chloride, HCl.

The solubility of HCl in water is 72.47g per 100g of water. As a percentage, this will be 72.47/(100+72.47)=0.42 or 42%.

The density of the HCl solution with a concentration of 42% is 1.208g/cm3 according to the reference data.

The molarity is calculated by the formula: cHCl=nHCl/Vsolution.

nHCl = mHCl/MHCl=72.47g/36.5 g*mol-1=1.985mol.

Vsolution=msolution/dsolution=(100 g+72.47 g)/(1.208 g/cm3)=142.77cm3=0.14277 l.

cHCl=1.985 mol/0.14277 l=13.9mol/l.

Answer: the molarity of hydrochloric acid is 13.9mol/l.